OpenStudy (anonymous):

Find F(1/3) if f(n) = 3n •1/3 •1 •3

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

Wherever you see an 'n', replace it with 1/3 Then evaluate and simplify

jimthompson5910 (jim_thompson5910):

\[\Large f(n) = 3n\] \[\Large f\left(\frac{1}{3}\right) = 3\left(\frac{1}{3}\right)\] \[\Large f\left(\frac{1}{3}\right) = \frac{3}{1}\left(\frac{1}{3}\right)\] \[\Large f\left(\frac{1}{3}\right) = \frac{3*1}{1*3}\] \[\Large f\left(\frac{1}{3}\right) = \frac{3}{3}\] \[\Large f\left(\frac{1}{3}\right) = 1\]